OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 9 interactions. If there are 4 children, there can be 18 interactions. Which recursive equation represents the pattern? (2 points)

OpenStudy (anonymous):

does anyone know the answer???

OpenStudy (anonymous):

let n = count of child, then: if 1 child a(1) = 0 if 2 child a(2) = 0 + 3 = a(1) + (2-1)*3 if 3 child a(3) = 3 + 6 = a(2) + (3-1)*3 if 4 child a(4) = 9 + 9 = a(3) + (4-1)*3 .... if n child a(n) = a(n-1) + (n-1)*3

OpenStudy (anonymous):

@ϟƘƦƖןןΣ✘

OpenStudy (anonymous):

so that is (n-1) cubed? if so thank you....it was the one question i could not figure out...

OpenStudy (psymon):

times 3

OpenStudy (anonymous):

times 3

OpenStudy (anonymous):

ohhh ok....thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you were correct